a train leaves a station and travels east at 40km/h three hours later a second train leaves on a parallel track traveling east at 60km/h how far away from the station do the trains meet?
Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

a train leaves a station and travels east at 40km/h three hours later a second train leaves on a parallel track traveling east at 60km/h how far away from the station do the trains meet?
------------------
1st train DATA:
rate = 40 km/hr ; time = x hrs ; distance = 40x km
-----
2nd train DATA:
rate = 60 km/hr ; time = x-3 hrs ; distance = 60(x-3) km
-----
Equation:
40x = 60x - 180
-20x = -180
x = 9 hrs (time for later train to catch up to earlier train)
------------
Cheers,
Stan H.

Answer by mananth(16949)

the distance traveled by both is same when one catches up with the other.
Let them meet at distance x
Train A speed= 40 km/h
Train B speed= 60 km/h

Time difference between the starting of the two= 3.00 hours

Time taken by Train A = X / 40
Time taken by Train B = X / 60

X / 40 - X / 60 = 3

LCD - 90
Multiply equation by 90 90

2.25 x - 1.5 x = 270
0.75 x = 270
/ 0.75
x= 360 km

Train B speed= 60 km/h
Distance = 360 miles
Time taken = 360 / 60
Time taken = 6 hours